What is the capacitance of a parallel plate capacitor with metal

plates, each of area 1.50 m2, separated by 2.01 mm? What charge is

stored in this capacitor if a voltage of 2.50 × 103 V is applied to it?


1
Expert's answer
2022-03-16T09:41:56-0400

The capacitance of the parallel plates capacitor


"C=\frac{\epsilon_0 A}{d}""C=\frac{8.85*10^{-12}*1.50}{2.01*10^{-3}}=6.60*10^{-9}\:\rm F"


The charge stored in capacitor

"q=CV\\
=6.60*10^{-9}*2.50*10^3=1.65*10^{-5}\:\rm C"